the length of a rectangle is five feet more than twice its width. the area is 133 sq ft. find dimensions of the rectangle.
SOMEBODY PLEASE HELP I KNOW I HAVE TO SET UP A COUPLE EQUATIONS BUT NOT SURE HOW. ANY HELP WOULD BE GREATLY APPRECIATED. THANKS IN ADVANCE.

Answer by stanbon(75887)

the length of a rectangle is five feet more than twice its width. the area is 133 sq ft. find dimensions of the rectangle.
------
Equations:
L = 2W+5
LW = 133
-----------
Substitute for "L" and solve for "W":
(2W+5)W = 133
---
2W^2+5W-133 = 0
---
Factor:
(W-7)(2W+19) = 0
---
Positive solution:
Width = 7 ft.
Solve for "Length":
L = 3W+5
L = 3*7+5
Length = 26 ft
================================
Cheers,
Stan H.